OpenStudy (anonymous):

Solve this system of equations using substitution, elimination or graphing. Show all of your work for the method you choose. y = 6x – 4 y = -2x + 28

OpenStudy (anonymous):

Well, we can see that y and y are very similar. Multiply a -1 to the top equation and then cancel out the y. That is your first step.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

You should have 0=-6x+4 0=-2x+28

OpenStudy (anonymous):

Now you can add them together to get 0=-8x+32

OpenStudy (anonymous):

Bring 32 over to the other side to get -32 = -8x. Can you solve x from here?

OpenStudy (anonymous):

do I divide -8 on both sides @KirbyLegs

OpenStudy (anonymous):

Yes you do. then the -8 will cancel out leaving you with 4 = x

OpenStudy (anonymous):

ok thats what I got! Thanks :)

OpenStudy (nirmalnema):

simply u can use double substitution method.. two different values of y are given. is you will equate it then you will get the value of x y = 6x – 4 y = -2x + 28 y=y 6x-4=-2x+28 transfer the like terms to one side 6x+2x=28+4 8x=32 x=4 put the value in any eq you will get y y=6(4)-4 =20 x=4 y=20

OpenStudy (anonymous):

ohh

OpenStudy (nirmalnema):

yup....

OpenStudy (anonymous):

thats interesting! thanks :)

OpenStudy (nirmalnema):

no problem thanks..